A 90% antifreeze solution is to be mixed with a 75% solution to make 120L of an 85% solution. How many liters of the 90% solution should be used?
Answer by rfer(16322)

0.90x+0.75(120-x)=0.85(120)
.90x+90-.75x=102
.15x=12
x=12/.15
x=80 lt. of 90% solution